# Higher-dimensional flying wing Steady Ricci Solitons

### v1

@@ Abstract.
For any $n\geq 4$, we construct an (n2)(n-2)-parameter family of steady gradient Ricci solitons with non-negative curvature operator and prescribed by the eigenvalues of Ricci tensor at the unique critical point of the soliton potential. Among them lies an (n3)(n-3)-parameter subfamily of non-collapsed solitons. These solitons generalized the flying wings constructed by the second named author.
Our approach is based on constructing continuous families of smooth Ricci flows emanating from continuous families of spherical polyhedra. This is built upon a combination of a new stability result of Ricci flows with scaling invariant estimates and the method of Gianniotis-Schulze in regularizing manifolds with singularities.
@@ 1.1. Flying wing steady solitons
In dimension two, Hamilton’s cigar soliton is the unique non-flat example [22]. In dimension three, the classical examples are the Bryant soliton and $\mathbb{R}\times\textnormal{Cigar}$. Hamilton further conjectured the existence of steady solitons asymptotic to sectors with angle $\alpha\in(0,\pi)$, known as flying wings. The second-named author confirmed this conjecture by constructing a one-paramter $\mathbb{Z}_{2}\times O(2)$-symmetric flying wings [28, 29]. Using the same method, she also constructed non-collapsed, $\mathbb{Z}_{2}\times O(n-1)$-symmetric flying wings for any dimension $n\geq 4$ [28]. Recently, Lavoyer-Peachey [31] generalized the method of the second named author to construct $O(p)\times O(q)$-symmetric $(p+q)$-dimensional flying wings for any pair of integers $p,q\geq 2$. See also for the works of Chan-Lai-Conlon [6] and Apostolov-Cifarelli [1] for Kähler flying wings.
In this paper, we widen the class of flying wings for $n\geq 4$ by constructing large families of steady solitons with non-negative curvature operator. In particular, we construct non-negatively curved steady solitons with prescribed Ricci curvature at the (unique) critical point and prescribed symmetric group:
@@
Among the steady solitons constructed in Theorem 1.1, it provides a large family of non-collapsed steady solitons.
@@
In addition to these non-collapsed families, it also provides new examples of collapsed $O(2)$-symmetric steady solitons.
Theorem 1.3.
For any $n\geq 3$, there exists an (n2)(n-2)-parameter family of pairwise different collapsed $O(2)$-symmetric steady gradient Ricci solitons with non-negative curvature operator.
@@
In comparison to mean curvature flow, the non-collapsed and collapsed steady gradient solitons in Theorem 1.2 and 1.3 are analogues of the mean curvature flow entire graphical translators and graphical translators on slabs of bounded width, which are constructed by Hoffman-Ilmanen-Martin-White [26].
In dimension three, the uniqueness of 3D flying wings remains open. In dimension four, it is conjectured by Haslhofer [25] that all $\kappa$-solutions (non-collapsed ancient flow with $\text{\rm Rm}\geq 0$) are the spherical solution, the cylindrical solution, Perelman’s oval, the 4D Bryant soliton, or the one-parameter family flying wings constructed by the second named author. Motivated by these, we conjecture that all non-collapsed steady gradient Ricci solitons with non-negative curvature operator and cylindrical tangent flow at infinity of the form $\mathbb{R}^{k}\times\mathbb{S}^{n-k}$, $n-k\geq 2$ is isometric (up to scaling) to one of the examples we constructed in Theorem 1.1.
@@
Consider the map from the expanding solitons to the Ricci eigenvalues at the critical point (note there are always two equal eigenvalues by the $O(2)$-symmetry), we obtain a smooth map from the simplex $\Omega$ to an (n2)(n-2)-dimensional simplex $\Delta$, with vertices the Ricci eigenvalues of the steady solitons corresponding to the vertices of $\Omega$. Then we show this map is surjective, thus by taking limits we obtain smooth steady solitons $\{\mathcal{S}_{\mathbf{x}}\}_{\mathbf{x}\in\Delta}$ parametrized by the simplex $\Delta$. The non-collapsed solitons in Theorem 1.2 is parametrized by a (n3)(n-3)-face of $\Delta$, and the collapsed solitons are in the complement.
@@ 1.2. Ricci flow starting from non-smooth initial data
so that $g_{0,(\beta_{2},\cdots,\beta_{n-1})}$ is the link of the spherical suspension and is the singular metrics on $\mathbb{S}^{n-2}$ with polyhedral singularity. In this work, we want to establish a canonical way to construct Ricci flow smoothing starting from these non-smooth initial metrics with desired properties.
We briefly discuss the idea behind and the subtly involved. By way of induction, if we have already constructed the Ricci flow smoothing $g_{(\beta_{2},\cdots,\beta_{n-1})}(t)$ coming out of the link $g_{0,(\beta_{2},\cdots,\beta_{n-1})}$, then we might consider the relatively less singular metric
@@
The challenging part is to show the continuity of $g_{(\beta_{1},\cdots,\beta_{n-1})}(t)$ as $(\beta_{1},\cdots,\beta_{n-1})$ varies. This is related to how one can gauge between $g_{(\beta_{1},\cdots,\beta_{n-1})}(t)$ and $g_{(\beta^{\prime}_{1},\cdots,\beta^{\prime}_{n-1})}(t)$ when they are initially close. When the initial data are smooth, it is very often to use the Ricci-harmonic map heat flow between two Ricci flows: $\partial_{t}F=\Delta_{g_{1}(t),g_{2}(t)}\,F,$ to transform a Ricci flow $g_{1}(t)$ to a Ricci-DeTurck flow $\hat{g}_{1}(t):=(F^{-1}(t))^{*}g_{1}(t)$ with respect to another Ricci flow $g_{2}(t)$, as long as $F(t)$ remains a diffeomorphism. The Ricci-DeTurck flow is a strictly parabolic system where the stability is relatively clear. In our case, it is tempting to mimic the smooth case by working on the regularized level, in the approximation scheme of Gianniotis-Schulze [20]. On the regularized level, the initial data is given by gluing in the expanding soliton constructed by Deruelle [17] near the singularities. And therefore in the asymptotic level, this is morally equivalent to ask if the Deruelle’s expander is continuously depending on the $L^{\infty}$ perturbation of its link. However, this $L^{\infty}$ stability of Deruelle’s expander is unclear in his construction [17], due to the absence of enough regularity on the link. We need to modify the construction, bypassing this issue.
The ultimate goal is to construct Ricci flows in such a way that the Ricci-harmonic map heat flow with certain continuity can be constructed. To achieve this, we construct Ricci flow $g_{(\beta_{1},\cdots,\beta_{n-1})}(t)$ coming out of $g_{0,(\beta_{1},\cdots,\beta_{n-1})}$ in a slightly different way so that the stability partially holds by construction. In short, the overall idea is to extend the existence using the above mentioned construction at one reference point $\hat{\beta}_{0}:=(\hat{\beta}_{1},\cdots,\hat{\beta}_{n-1})$ in the parameter space and propagate the existence to other $\beta$} using weak stability. More precisely, we fix a reference point $\hat{\beta}_{0}$ and implement the above mentioned strategy to obtain a reference Ricci flow $g_{(\hat{\beta}_{1},\cdots,\hat{\beta}_{n-1})}(t)$. Instead of repeating the construction using the same method on a different $\beta$}, we consider construction using weak stability with respect to $g_{(\hat{\beta}_{1},\cdots,\hat{\beta}_{n-1})}(t)$.
@@
Indeed, the metrics are allowed to be non-smooth initially to some extent, which will be important for applications. We refer readers to Section 4 for a more general statement of the stability result which allows $L^{\infty}$ type singularity on the initial data.
@@
Arguing in this way inductively, we construct Ricci flow $g_{{\beta},\tau}(t)$ for all $\beta$} in the parameter space. In practice, we first carried this out in the $\tau$-level, i.e. we will consider the stability of Ricci-DeTurck flow from $g_{0,(\beta_{1},\cdots,\beta_{n-1}),\tau}$, and then let $\tau\to 0$. The gauge fixing using harmonic map heat flow in the earlier dimension will be involved in order to compare different regularized links $g_{(\beta_{2},\cdots,\beta_{n-1})}(\tau)$ in $L^{\infty}$ topology. Similarly whenever a different reference Ricci flow is used, we need to re-fix the gauge of the initial data due to the Ricci-DeTurck ODE in earlier step and gauge fixing in earlier dimension, so that stability with respect to the new Ricci flow applies in the new gauge. Despite the technicality, the stability of the new construction can be achieved easily for $\beta$} and $\hat{\beta}$ living along the same ray. The remaining subtly lies in the stability when $\beta$} and $\hat{\beta}$ are close to each other but not necessarily along the same ray. Particularly, the Ricci flows are constructed under a different line of background Ricci flows. We manage to overcome this by tracing the gauge change carefully, and show that Ricci-harmonic map heat flow between $g_{\beta}(t)$ and $g_{\hat{\beta}}(t)$ can still be solved with stability estimates, whenever $^{\beta},\hat{\beta}$ are close. Equivalently, $g_{\beta}(t)$ can be transformed into a Ricci-DeTurck flow with respect to $g_{\hat{\beta}}(t)$ through a gauge change using Ricci-harmonic map heat flow. This will be achieved by careful approximation procedure and will be discussed in detail in Section 5.
@@ Acknowledgment
The authors would like to thank Felix Schulze for discussion related to stability of Ricci flow. P.-Y. Chan is supported by the Yushan Young Fellow Program of the Ministry of Education (MOE), Taiwan (MOE-108-YSFMS-0004-012-P1), and by the NSTC grant 113-2115-M-007 -014 -MY2. Y. Lai is supported by NSF grant DMS-2506832. M.-C. Lee is supported by Hong Kong RGC grant (Early Career Scheme) of Hong Kong No. 24304222 and No. 14300623, NSFC grant No. 12222122 and an Asian Young Scientist Fellowship.
@@ 3.1. Expander with $\mathrm{Rm}\geq 0$
Then $g_{N}(t):=t(\Phi_{t})^{*}g_{N,1}$ defines a Ricci flow on $N\times(0,+\infty)$. Moreover it follows from Hamilton-Perelman distance distortion, for instances see [43, Lemma 3.1], that $d_{g_{N}(t)}$ converges to some distance function $d_{0}$ on $N$ as $t\to 0$ so that $(N,d_{0},p_{N})$ is isometric to $C(X)$ as metric space. ∎
@@
Since $g_{N}(0)=g_{c,X}=dr^{2}+r^{2}g_{X}$ and $X:=(\mathbb{S}^{n-1},h)$ is smooth, the conclusion follows from [12, Corollary 3.2] and the modified Shi’s type estimates [13, Theorem 14.16], see also [39, Theorem 10.3]. ∎
@@ 3.2. Existence of Ricci flow
for $s$ sufficiently small. Here we have used [43, Lemma 3.1] and Lemma 3.4. In particular, $g_{s,0}=g_{N}(s)$ on $B_{c}(o,s^{1/4})$ so that $\mathrm{Rm}(g_{s,0})>0$. This proves the estimate on Case 1. If $x\notin\overline{B_{c}(o,3s^{1/4})}$, then for $s$ sufficiently small,
@@
Here we have used [43, Lemma 3.1] and Lemma 3.4. Thus, $g_{s,0}=g_{0}$ where we have $\mathrm{Rm}_{-}(g_{s,0})\leq 1$. This proves the estimate on Case 3.
@@
It follows from the proof of [5, Theorem 4.3] and Proposition 3.6 that $(M_{s},g_{s,0})\in\mathcal{K}_{IC1}(n,f,v)$ in the sense of [33, Definition 1.2], uniformly for $f(t)=C_{2}(n,Z)\cdot t^{\sigma/2}$ and some $v>0$, as $s\to 0$. Hence, [33, Theorem 1.1] applies to obtain a one parameter family of Ricci flow $g_{s}(t),t\in[0,S]$ starting from $g_{s,0}$ satisfying
@@
on $(0,S]$, for some uniform $v_{0},{\alpha},S>0$ independent of $s\to 0$. Here conclusion (c) is from [33, Proposition 3.2 & Lemma 3.1]. This enables us to take sub-sequential limit of $g_{s_{i}}(t)$ as $s_{i}\to 0$, by Hamilton’s compactness [23] to obtain a Ricci flow $(M,g(t)),t\in(0,S]$. By Proposition 3.6, [33, Lemma 4.1] and estimates above, $(M,g(t))$ converges to $(Z,g_{0})$ as $t\to 0$ in the Gromov-Hausdorff sense. By [19, Theorem 1.6] or local regularity away from tips, $g(t)$ converges smoothly to some smooth metric $g(0)$ as $t\to 0$ away from $\{p_{i}\}_{i=1}^{N}$. Moreover, $d_{g_{0}}$ is distance-isometric to $d_{g(0)}$ via a Lipschitz map $\Psi:Z\to M$. By Myers–Steenrod, $\Psi$ is smooth outside $\{p_{i}\}_{i=1}^{N}$. We might assume $\Psi=\mathrm{Id}$ for notation convenience and $d_{g(t)}\to d_{0}=d_{g_{0}}$.
@@
Since $r\in(0,1]$, we might assume $r=1$ by considering $r^{-2}g(r^{2}t)$. Since $\varphi(0)\leq\varepsilon^{2}r^{2}$ on $B_{g_{0}}(x,2)$, we might apply [32, Theorem 1.1] so that for all $\ell>0$, there is $\hat{T}_{\ell}>0$ such that for all $t\in[0,\hat{T}_{\ell}$],
@@
then it satisfies $v\geq 0$ by $\mathrm{Rm}\geq 0$, and $(\partial_{t}-\Delta_{g(t)})v\geq 0$ in the sense of barrier by [21]. We follow the idea in the proof of [6, Lemma 4.2]. Let $K(x,t;y,s)$ be the heat kernel for $\partial_{t}-\Delta_{g(t)}$. From smooth convergence away from $p$, for any $\delta,\varepsilon>0$, there is $s_{0}>0$ such that $v\geq 1-\varepsilon$ on $B_{g_{0}}(p,\delta)^{c}\times(0,s_{0})$. By the kernel representation, for all $t\in(0,S]$,

### v2

@@ Abstract.
For any $n\geq 4$, we construct an (n2)(n-2)-parameter family of steady gradient Ricci solitons with non-negative curvature operator and prescribed by the eigenvalues of Ricci tensor at a critical point of the soliton potential. Among them lies an (n3)(n-3)-parameter subfamily of non-collapsed solitons. These solitons generalized the flying wings constructed by the second named author and produced new examples of steady gradient Ricci solitons with non-negative curvature operator for $n\geq 4$. Our approach is based on constructing continuous families of Ricci flows smoothing emanating from continuous families of spherical polyhedra which still preserves symmetry.
This is built upon a new stability result of Ricci flows with scaling invariant estimates. As another application of the method, we prove the stability of asymptotically conical expanding solitons constructed by Deruelle under $L^{\infty}$ perturbation of links. In particular, the $C^{0}$-convergence of smooth links implies the smooth convergence of the expanding solitons.
@@ 1.1. Flying wing steady solitons
In dimension two, Hamilton’s cigar soliton is the unique non-flat example [24]. In dimension three, the classical examples are the Bryant soliton and $\mathbb{R}\times\textnormal{Cigar}$. Hamilton further conjectured the existence of steady solitons asymptotic to sectors with angle $\alpha\in(0,\pi)$, known as flying wings. The second-named author confirmed this conjecture by constructing a one-paramter $\mathbb{Z}_{2}\times O(2)$-symmetric flying wings [30, 31]. Using the same method, she also constructed non-collapsed, $\mathbb{Z}_{2}\times O(n-1)$-symmetric flying wings for any dimension $n\geq 4$ [30]. Recently, Lavoyer-Peachey [33] generalized the method of the second named author to construct $O(p)\times O(q)$-symmetric $(p+q)$-dimensional flying wings for any pair of integers $p,q\geq 2$. See also for the works of Chan-Conlon-Lai [8] and Apostolov-Cifarelli [1] for Kähler flying wings.
In this paper, we widen the class of flying wings for $n\geq 4$ by constructing large families of steady solitons with non-negative curvature operator. In particular, we construct non-negatively curved steady solitons with prescribed Ricci curvature at a critical point of $f$ and prescribed symmetric group:
@@
Among the steady solitons constructed in Theorem 1.1, it provides a large family of non-collapsed steady solitons, using a similar proof as [30, Corollary 3.5].
@@
In previous works [1, 4, 6, 8, 30, 31, 33], all known steady gradient Ricci solitons of positive sectional curvature have at most $2$ distinct eigenvalues of Ric at the unique critical point of $f$. With a more judicious choice of $\lambda_{i}$, Theorem 1.1 yields large families of new examples of steady solitons with positive curvature operator and at least $3$ distinct eigenvalues of Ric at the unique critical point of $f$ in dimensions four and higher, see Remark 6.1.
In comparison to mean curvature flow, the non-collapsed steady gradient solitons in Theorem 1.1 is analogues of the mean curvature flow entire graphical translators and graphical translators on slabs of bounded width, which are constructed by Hoffman-Ilmanen-Martin-White [28].
In dimension three, the uniqueness of 3D flying wings remains open. In dimension four, it is conjectured by Haslhofer [27] that all $\kappa$-solutions (non-collapsed ancient flow with $\text{\rm Rm}\geq 0$) are the spherical solution, the cylindrical solution, Perelman’s oval, $\mathbb{CP}^{2}$, the 4D Bryant soliton, the one-parameter family flying wings constructed by the second named author, or the one-parameter ovals constructed by Haslhofer [27]. Motivated by these, we conjecture that all non-collapsed steady gradient Ricci solitons with non-negative curvature operator and cylindrical tangent flow at infinity of the form $\mathbb{R}^{k}\times\mathbb{S}^{n-k}$, $n-k\geq 2$ is isometric (up to scaling) to one of the examples we constructed in Theorem 1.1.
@@
Consider the map from the expanding solitons to the Ricci eigenvalues at the critical point (note there are always two equal eigenvalues by the $O(2)$-symmetry), we obtain a smooth map from the simplex $\Omega$ to an (n2)(n-2)-dimensional simplex $\Delta$, with vertices the Ricci eigenvalues of the steady solitons corresponding to the vertices of $\Omega$. Then we show this map is surjective, thus by taking limits we obtain smooth steady solitons $\{\mathcal{S}_{\mathbf{x}}\}_{\mathbf{x}\in\Delta}$ parametrized by the simplex $\Delta$. The non-collapsed solitons in Theorem 1.2 is parametrized by a (n3)(n-3)-face of $\Delta$.
@@ 1.2. Ricci flow starting from non-smooth initial data
so that $g_{0,(\beta_{2},\cdots,\beta_{n-1})}$ is the link of the spherical suspension and is the singular metrics on $\mathbb{S}^{n-2}$ with polyhedral singularity. In this work, we want to establish a canonical way to construct continuous Ricci flow smoothing starting from these non-smooth initial metrics with desired properties as $\beta$} varies, see Theorem 5.2 for the detailed statement.
We briefly discuss the idea behind and the subtlety involved. By way of induction, if we have already constructed the Ricci flow smoothing $g_{(\beta_{2},\cdots,\beta_{n-1})}(t)$ coming out of the link $g_{0,(\beta_{2},\cdots,\beta_{n-1})}$, then we might consider the relatively less singular metric
@@
The challenging part is to show the continuity of $g_{(\beta_{1},\cdots,\beta_{n-1})}(t)$ as $(\beta_{1},\cdots,\beta_{n-1})$ varies. This is related to how one can gauge between $g_{(\beta_{1},\cdots,\beta_{n-1})}(t)$ and $g_{(\beta^{\prime}_{1},\cdots,\beta^{\prime}_{n-1})}(t)$ when they are initially close. When the initial data are smooth, it is very often to use the Ricci-harmonic map heat flow between two Ricci flows: $\partial_{t}F=\Delta_{g_{1}(t),g_{2}(t)}\,F,$ to transform a Ricci flow $g_{1}(t)$ to a Ricci-DeTurck flow $\hat{g}_{1}(t):=(F^{-1}(t))^{*}g_{1}(t)$ with respect to another Ricci flow $g_{2}(t)$, as long as $F(t)$ remains a diffeomorphism. The Ricci-DeTurck flow is a strictly parabolic system where the stability is relatively clear. In our case, it is tempting to mimic the smooth case by working on the regularized level, in the approximation scheme of Gianniotis-Schulze [22]. On the regularized level, the initial data is given by gluing in the expanding soliton constructed by Deruelle [19] near the singularities and thus the tangent flow of the Ricci flow is governed by the expanding Ricci soliton constructed by Deruelle [19]. Deruelle’s argument is based on PDE deformation and continuity method. In particular, the stability of those expanders are only known to be true as the link varies in $C^{3,{\alpha}}$ topology, while our link of polyhedral singularties varies only in $L^{\infty}$ topology. Therefore, we need to modify the gluing construction, compatible with $L^{\infty}$ variation of links.
The ultimate goal is to construct Ricci flows so that the Ricci flow with singular initial data can be gauged with each other whenever the initial singular metrics are close in weak sense, i.e. in $L^{\infty}$ sense. To achieve this, we construct Ricci flow $g_{(\beta_{1},\cdots,\beta_{n-1})}(t)$ coming out of $g_{0,(\beta_{1},\cdots,\beta_{n-1})}$ in a slightly different way so that the stability partially holds by construction. In short, the overall idea is to extend the existence using the above mentioned construction at one reference point $\hat{\beta}_{0}:=(\hat{\beta}_{1},\cdots,\hat{\beta}_{n-1})$ in the parameter space and propagate the existence to other $\beta$} using weak stability. More precisely, we fix a reference point $\hat{\beta}_{0}$ and implement the above mentioned strategy to obtain a reference Ricci flow $g_{(\hat{\beta}_{1},\cdots,\hat{\beta}_{n-1})}(t)$. Instead of repeating the construction using the same method on a different $\beta$}, we consider construction using weak stability with respect to $g_{(\hat{\beta}_{1},\cdots,\hat{\beta}_{n-1})}(t)$.
@@
In particular, this generalizes the stability in [5, 20, 29] where the reference Ricci flow $\tilde{g}(t)$ is assumed to have bounded curvature. Indeed, the metrics are allowed to be non-smooth initially to some extent, which will be important for applications. We refer readers to Section 4 for a more general statement of the stability result which allows $L^{\infty}$ type singularity on the initial data.
@@
Arguing in this way inductively, we construct Ricci flow $g_{{\beta},\tau}(t)$ for all $\beta$} in the parameter space. In practice, we first carry this out in the $\tau$-level, i.e. we will consider the stability of Ricci-DeTurck flow from $g_{0,(\beta_{1},\cdots,\beta_{n-1}),\tau}$, and then let $\tau\to 0$. The gauge fixing using harmonic map heat flow in the earlier dimension will be involved in order to compare different regularized links $g_{(\beta_{2},\cdots,\beta_{n-1})}(\tau)$ in $L^{\infty}$ topology. Similarly whenever a different reference Ricci flow is used, we need to re-fix the gauge of the initial data due to the Ricci-DeTurck ODE in earlier step and gauge fixing in earlier dimension, so that stability with respect to the new Ricci flow applies in the new gauge. Despite the technicality, the stability of the new construction can be achieved easily for $\beta$} and $\hat{\beta}$ living along the same ray. The remaining subtlety lies in the stability when $\beta$} and $\hat{\beta}$ are close to each other but not necessarily along the same ray. Particularly, the Ricci flows are constructed under a different line of background Ricci flows. We manage to overcome this by tracing the gauge change carefully, and show that Ricci-harmonic map heat flow between $g_{\beta}(t)$ and $g_{\hat{\beta}}(t)$ can still be solved with stability estimates, whenever $^{\beta},\hat{\beta}$ are close. Equivalently, $g_{\beta}(t)$ can be transformed into a Ricci-DeTurck flow with respect to $g_{\hat{\beta}}(t)$ through a gauge change using Ricci-harmonic map heat flow. This will be achieved by careful approximation procedure and will be discussed in detail in Section 5.
The new construction of the Ricci flow smoothing is possibly different from the flow constructed by Gianniotis-Schulze [22], even in case of isolated singularities. The uniqueness and stability of Ricci flow starting from singular initial metric are still unclear in general. Restricting to the case when the flow is coming out of a cone with smooth link with $\mathrm{Rm}>1$, it was shown by Deruelle [19] that the expanding soliton with positive curvature coming of that cone is unique. On the other hand, the Ricci flow smoothing produced by the new construction scheme can be shown to be an expander with positive curvature, using the result in [10]. Consequently, two solutions will coincide with each other.
In particular, thanks to the nature of new construction, this shows that Deruelle’s expanders are stable with respect to $L^{\infty}$ perturbation of the (smooth) link. As a corollary of the stability, we obtain that the $C^{0}$-convergence of links implies the $C^{\infty}$-convergence of the expanding solitons.
Theorem 1.4.
Let $\{(\mathbb{S}^{n-1},h_{i})\}_{i=1}^{\infty}$ and $(\mathbb{S}^{n-1},h_{\infty})$ be a sequence of smooth Riemannian manifolds. Suppose $\text{\rm Rm}(h_{i})>1$, for all $i\in\mathbb{N}\cup\{\infty\}$, and
$\|h_{i}-h_{\infty}\|_{C^{0}(\mathbb{S}^{n-1})}\to 0,\qquad\textnormal{as }i\to\infty.$
Let $\{(M_{i},g_{i},f_{i},p_{i})\}_{i=1}^{\infty}$ be the unique Ricci solitons coming out of the cone over $(\mathbb{S}^{n-1},h_{i})$, constructed by Deruelle [19]. Then there are smooth diffeomorphisms $\Phi_{i}$ on $\mathbb{S}^{n-1}$ such that for any $k\in\mathbb{N}$,
$\|\Phi_{i}^{*}g_{i}-g_{\infty}\|_{C^{k}(\mathbb{S}^{n-1})}\to 0,\qquad\textnormal{as }i\to\infty.$
We refer readers to Theorem 5.4 for the quantitative version of the expander’s stability.
@@ Acknowledgment
The authors would like to thank Felix Schulze for fruitful discussion related to stability of Ricci flow. P.-Y. Chan is supported by the Yushan Young Fellow Program of the Ministry of Education (MOE), Taiwan (MOE-108-YSFMS-0004-012-P1), and by the NSTC grant 113-2115-M-007 -014 -MY2. Y. Lai is supported by NSF grant DMS-2506832. M.-C. Lee is supported by Hong Kong RGC grant (Early Career Scheme) of Hong Kong No. 24304222, No. 14300623 and No. 14304225, NSFC grant No. 12222122 and an Asian Young Scientist Fellowship.
@@ 3.1. Expander with $\mathrm{Rm}\geq 0$
Then $g_{N}(t):=t(\Phi_{t})^{*}g_{N,1}$ defines a Ricci flow on $N\times(0,+\infty)$. Moreover it follows from Hamilton-Perelman distance distortion, for instances see [45, Lemma 3.1], that $d_{g_{N}(t)}$ converges to some distance function $d_{0}$ on $N$ as $t\to 0$ so that $(N,d_{0},p_{N})$ is isometric to $C(X)$ as metric space. ∎
@@
Since $g_{N}(0)=g_{c,X}=dr^{2}+r^{2}g_{X}$ and $X:=(\mathbb{S}^{n-1},h)$ is smooth, the conclusion follows from [14, Corollary 3.2] and the modified Shi’s type estimates [15, Theorem 14.16], see also [41, Theorem 10.3]. ∎
@@ 3.2. Existence of Ricci flow
for $s$ sufficiently small. Here we have used [45, Lemma 3.1] and Lemma 3.4. In particular, $g_{s,0}=g_{N}(s)$ on $B_{c}(o,s^{1/4})$ so that $\mathrm{Rm}(g_{s,0})>0$. This proves the estimate on Case 1. If $x\notin\overline{B_{c}(o,3s^{1/4})}$, then for $s$ sufficiently small,
@@
Here we have used [45, Lemma 3.1] and Lemma 3.4. Thus, $g_{s,0}=g_{0}$ where we have $\mathrm{Rm}_{-}(g_{s,0})\leq 1$. This proves the estimate on Case 3.
@@
It follows from the proof of [7, Theorem 4.3] and Proposition 3.6 that $(M_{s},g_{s,0})\in\mathcal{K}_{IC1}(n,f,v)$ in the sense of [35, Definition 1.2], uniformly for $f(t)=C_{2}(n,Z)\cdot t^{\sigma/2}$ and some $v>0$, as $s\to 0$. Hence, [35, Theorem 1.1] applies to obtain a one parameter family of Ricci flow $g_{s}(t),t\in[0,S]$ starting from $g_{s,0}$ satisfying
@@
on $(0,S]$, for some uniform $v_{0},{\alpha},S>0$ independent of $s\to 0$. Here conclusion (c) is from [35, Proposition 3.2 & Lemma 3.1]. This enables us to take sub-sequential limit of $g_{s_{i}}(t)$ as $s_{i}\to 0$, by Hamilton’s compactness [25] to obtain a Ricci flow $(M,g(t)),t\in(0,S]$. By Proposition 3.6, [35, Lemma 4.1] and estimates above, $(M,g(t))$ converges to $(Z,g_{0})$ as $t\to 0$ in the Gromov-Hausdorff sense. By [21, Theorem 1.6] or local regularity away from tips, $g(t)$ converges smoothly to some smooth metric $g(0)$ as $t\to 0$ away from $\{p_{i}\}_{i=1}^{N}$. Moreover, $d_{g_{0}}$ is distance-isometric to $d_{g(0)}$ via a Lipschitz map $\Psi:Z\to M$. By Myers–Steenrod, $\Psi$ is smooth outside $\{p_{i}\}_{i=1}^{N}$. We might assume $\Psi=\mathrm{Id}$ for notation convenience and $d_{g(t)}\to d_{0}=d_{g_{0}}$.
@@
Since $r\in(0,1]$, we might assume $r=1$ by considering $r^{-2}g(r^{2}t)$. Since $\varphi(0)\leq\varepsilon^{2}r^{2}$ on $B_{g_{0}}(x,2)$, we might apply [34, Theorem 1.1] so that for all $\ell>0$, there is $\hat{T}_{\ell}>0$ such that for all $t\in[0,\hat{T}_{\ell}$],
@@
then it satisfies $v\geq 0$ by $\mathrm{Rm}\geq 0$, and $(\partial_{t}-\Delta_{g(t)})v\geq 0$ in the sense of barrier by [23]. We follow the idea in the proof of [8, Lemma 4.2]. Let $K(x,t;y,s)$ be the heat kernel for $\partial_{t}-\Delta_{g(t)}$. From smooth convergence away from $p$, for any $\delta,\varepsilon>0$, there is $s_{0}>0$ such that $v\geq 1-\varepsilon$ on $B_{g_{0}}(p,\delta)^{c}\times(0,s_{0})$. By the kernel representation, for all $t\in(0,S]$,